a ball moving with a velocity 20m/s has a mass of 50gm. if collides against a wall normally and is rebounded normally with the same speed. if the time of impact of the ball and the wall is 40 mili seconds, the average force exerted on wall in dunes is

Grade:11

1 Answers

Karan Gupta
6 years ago
Change in velocity = 20 m/s – (-20 m/s) = 40 m/s
Change in momentum of ball = 0.05 * 40 = 2 kg-m/s
Average force = 2 / (40 * 0.001) = 50 N = 5 mega dynes